# Uncertainty Quantification for Gradient-based Explanations in Neural Networks

### v1

@@ Abstract
Keywords:
uncertainty estimation methods, explanation methods, explanation distribution, coefficient of variation, pixel insertion/deletion
@@ 1 Introduction
(a) Input
Image
(b) Explanation
Mean
(c) Explanation
Uncertainty
Figure 1: Visualization of explanation with uncertainty on an image from FER+ dataset using Guided Backpropagation as explanation method and Deep Ensembles as uncertainty estimation method.
@@ 3.1 Explanation Uncertainty
$\mu(x)=T^{-1}\sum_{i}f_{\theta}(x)$ (7)
$\sigma^{2}(x)=T^{-1}\sum_{i}(f_{\theta}(x)-\mu(x))^{2}$ (8)
@@
where $\text{E}_{\sigma}(x)$ and $\text{E}_{\mu}(x)$ are the standard deviation and mean explanations respectively. Note that the explanation uncertainty is explicitly concerned with modeling uncertainty in the explanation and is unrelated to uncertainty explanation where the objective is to explain the uncertainty in a given prediction.
@@ 4 Experimental Setup
4.1 Explanation Uncertainty for Image Classification
We test the proposed approach for the image classification task using the CIFAR-10 and the FER+ dataset. For this, we use the {Deep Ensemble, MC-Dropout, MC-DropConnect, Flipout} as uncertainty estimation methods and {Guided Backpropagation, Integrated Gradients} as the explanation methods. This provides us with a total of eight combinations of uncertainty estimation and explanation methods that we test. It should be noted that the LIME has not been used along with the image classification task as it is computationally expensive.
@@ 4.1 Explanation Uncertainty for Image Classification
4.2 Pixel Deletion/Insertion
We focus on quantifying the quality of the generated explanations using our proposed approach. To this end, we identify the pixel deletion and the pixel insertion metric proposed by [10] in their work. The objective is to compute the changes in the predicted class score by the subsequent addition/removal of pixels to/from an image. We plot this change in the class score against the number of pixels added/deleted and compute the area under this curve (AUC) which is the pixel insertion/deletion metric. The addition/removal of the pixels to/from the input image is governed by the explanation heatmap associated with the image under consideration.
@@ 4.2 Pixel Deletion/Insertion
4.3 Explanation Uncertainty - Prediction vs Ground Truth Neurons
We propose modifying the computation of explanation uncertainty by varying the gradient computation step in our original approach. This modification is only tested with the FER+ dataset and is explained with the help of Figure 3.
Figure 3: Illustration to describe the different approaches to compute gradients and generate explanations. The network predicted the expression in the input image to be happy whereas the ground truth expression is suprise.
In our proposed approach, the explanation is generated by calculating the gradient of the network output with respect to the input image. Mathematically, this is formulated as:
$E(x)=\frac{\partial(\text{Output}_{\text{Prediction}})}{\partial(\text{Input}\ \text{Image})}$ (13)
where $Output_{Prediction}$ is the same as network prediction from Figure 3. The explanation generated from this calculation helps us understand why the network predicted what it predicted.
However, now we modify the gradient computation step as:
$E(x)=\frac{\partial(\text{Output}_{\text{Ground\ Truth}})}{\partial(\text{Input\ Image)}}$ (14)
where $\text{Output}_{\text{Ground\ Truth}}$ is the activation of the unit corresponding to the ground truth (GT) as depicted in Figure 3. The explanation generated using this calculation helps to visualize the features/image regions that the network should have focused on for its prediction to match the ground truth. It is important to note that Equation 13 and Equation 14 converge when the network prediction matches the ground truth for a given input.
@@ 4.3 Explanation Uncertainty - Prediction vs Ground Truth Neurons
4.4 Explanation Uncertainty for Tabular Regression
@@ 5 Experimental Results
In this section, we present the results of the experiments discussed in the previous section.
5.1 Explanation Uncertainty for Image Classification
The explanation heatmaps for the CIFAR-10 and the FER+ datasets have been visualized in Figure 4 and Figure 5 respectively. As previously described, the size of these heatmaps is the same as the size of the input image. In the subsequent discussions, the heatmaps will be referred to by their identifier namely the mean explanation heatmap as $\mu$, the standard deviation explanation heatmap as $\sigma$, and the coefficient of variation explanation heatmap as $CV$.
@@ 5.1 Explanation Uncertainty for Image Classification
Figure 5 shows the input image with the expression suprise taken from the FER+. It can be observed in the $\mu$ heatmaps that the GBP highlights the relevant facial features to a greater extent as compared to IG. This observation is supported by the $\sigma$ heatmaps of both these methods as well. The $\sigma$ and the $CV$ heatmaps generated by the GBP have low activation around the relevant pixel regions for all the uncertainty estimation methods. This implies that the explanations generated by the GBP have a higher agreement and are therefore highly confident. However, this is not the case for IG where the pixel activations in the $\mu$ heatmaps are scattered. This indicates a low confidence in the generated explanation.
@@
$\mu$-IG $\sigma$-IG CV-IG $\mu$-GBP $\sigma$-GBP CV-GBP
Refer to caption Deep Ensemble Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-Dropout Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-DropConnect Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: The visualization of the mean ($\mu$), the standard deviation ($\sigma$), and the coefficient of variation ($CV$) heatmaps generated using the combinations of uncertainty estimation methods (Deep Ensemble, MC-Dropout, MC-DropConnect and Flipout) and explanation methods (Guided Backpropagation and Integrated Gradients). The uncertainty estimation methods are organized in individual rows and the explanation methods are aligned along the columns with the first three columns representing the $\mu$, $\sigma$, and the $CV$ heatmaps for IG and the last three columns representing the same for GBP. The image belongs to the CIFAR-10 dataset.
$\mu$-IG $\sigma$-IG CV-IG $\mu$-GBP $\sigma$-GBP CV-GBP
Refer to caption Deep Ensemble Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-Dropout Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-DropConnect Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: The visualization of the mean ($\mu$), the standard deviation ($\sigma$), and the coefficient of variation ($CV$) heatmaps generated using the combinations of uncertainty estimation methods (Deep Ensemble, MC-Dropout, MC-DropConnect and Flipout) and explanation methods (Guided Backpropagation and Integrated Gradients). The uncertainty estimation methods are organized in individual rows and the explanation methods are aligned along the columns with the first three columns representing the $\mu$, $\sigma$, and the $CV$ heatmaps for IG and the last three columns representing the same for GBP. The image belongs to the FER+ dataset.
5.2 Pixel Deletion/Insertion
The results for pixel deletion/insertion are provided in Figures 6 - Figure 9 and Table 1 and Table 2. The figures depict the variation in the class score for a batch of images (from both the CIFAR-10 and FER+) by applying the pixel insertion/deletion based on their corresponding explanation heatmaps ($\mu$ and $\sigma$). The tables contain the AUCs for these plots. It should be noted that a higher AUC for pixel insertion and a lower AUC for pixel deletion is desired to ensure good-quality explanations. The pixel insertion and deletion curves have been plotted for batches of images representing all the classes in both datasets (CIFAR-10 and FER+). The order of pixel insertion/deletion has been determined using the mean and the standard deviation explanation heatmaps of the corresponding images from these datasets.
It is evident from these figures and tables that for CIFAR-10 the best explanations are generally generated for the frog class and the worst explanations are generated for the dog class. These observations are supported by high AUC for pixel insertion and low AUC for pixel deletion for these classes.
@@ 5.2 Pixel Deletion/Insertion
Deletion
$\mu$-GBP
$\sigma$-GBP Refer to caption
$\mu$-IG Refer to caption
$\sigma$-IG Refer to caption
Insertion
Deletion
Insertion
Figure 6: Pixel flipping curves for Deep Ensemble (rows 1-2) and MC-Dropout (rows 3-4) for CIFAR-10. The first and the third rows depict the deletion curves whereas the second and the fourth rows depict the insertion curves. From left to right, the columns show the following heatmaps and the explanation methods they correspond to: (i) $\mu$-GBP (ii) $\sigma$-GBP (iii) $\mu$-IG (iv) $\sigma$-IG.
Deletion
$\mu$-GBP
$\sigma$-GBP
$\mu$-IG
$\sigma$-IG
Insertion
Deletion
Insertion
Figure 7: Pixel flipping curves for MC-DropConnect (rows 1-2) and Flipout (rows 3-4) for CIFAR-10. The first and the third rows depict the deletion curves whereas the second and the fourth rows depict the insertion curves. From left to right, the columns show the following heatmaps and the explanation methods they correspond to: (i) $\mu$-GBP (ii) $\sigma$-GBP (iii) $\mu$-IG (iv) $\sigma$-IG.
truck frog bird cat deer plane ship auto dog horse
$\mu$-GBP-E 0.80 0.62 0.86 0.65 0.48 0.52 0.40 0.47 0.47 0.52 0.34 0.69 0.61 0.53 0.59 0.57 0.30 0.41 0.52 0.56
$\sigma$-GBP-E 0.78 0.59 0.87 0.69 0.47 0.51 0.40 0.45 0.45 0.50 0.34 0.66 0.62 0.51 0.59 0.56 0.31 0.39 0.51 0.54
$\mu$-IG-E 0.79 0.72 0.86 0.73 0.48 0.53 0.40 0.47 0.46 0.56 0.43 0.62 0.58 0.62 0.57 0.73 0.32 0.43 0.58 0.65
$\sigma$-IG-E 0.78 0.66 0.86 0.77 0.48 0.51 0.39 0.45 0.48 0.54 0.40 0.54 0.56 0.59 0.57 0.68 0.32 0.41 0.58 0.56
$\mu$-GBP-O 0.69 0.29 0.86 0.39 0.56 0.39 0.45 0.25 0.53 0.29 0.37 0.45 0.65 0.33 0.68 0.38 0.30 0.19 0.42 0.24
$\sigma$-GBP-O 0.70 0.29 0.85 0.40 0.53 0.40 0.44 0.27 0.52 0.32 0.38 0.41 0.62 0.34 0.66 0.38 0.28 0.19 0.43 0.23
$\mu$-IG-O 0.75 0.38 0.82 0.48 0.54 0.39 0.43 0.28 0.61 0.35 0.47 0.38 0.56 0.40 0.69 0.49 0.39 0.19 0.56 0.34
$\sigma$-IG-O 0.73 0.36 0.83 0.50 0.59 0.37 0.39 0.27 0.60 0.36 0.47 0.32 0.57 0.41 0.66 0.47 0.36 0.17 0.51 0.31
$\mu$-GBP-C 0.71 0.29 0.85 0.36 0.47 0.29 0.50 0.29 0.64 0.25 0.36 0.51 0.69 0.25 0.70 0.29 0.41 0.33 0.52 0.24
$\sigma$-GBP-C 0.69 0.30 0.85 0.42 0.45 0.29 0.45 0.30 0.63 0.29 0.33 0.44 0.69 0.25 0.67 0.28 0.37 0.30 0.49 0.23
$\mu$-IG-C 0.72 0.38 0.85 0.42 0.50 0.25 0.46 0.32 0.62 0.33 0.46 0.41 0.63 0.33 0.64 0.49 0.43 0.37 0.63 0.35
$\sigma$-IG-C 0.72 0.34 0.83 0.46 0.43 0.25 0.38 0.32 0.65 0.36 0.45 0.30 0.62 0.31 0.65 0.47 0.42 0.31 0.60 0.33
$\mu$-GBP-F 0.70 0.27 0.92 0.39 0.59 0.38 0.51 0.33 0.51 0.22 0.45 0.39 0.59 0.20 0.73 0.35 0.31 0.23 0.55 0.25
$\sigma$-GBP-F 0.61 0.26 0.92 0.47 0.57 0.40 0.44 0.34 0.51 0.23 0.41 0.32 0.59 0.20 0.69 0.33 0.29 0.22 0.54 0.22
$\mu$-IG-F 0.70 0.36 0.89 0.49 0.61 0.42 0.48 0.36 0.58 0.29 0.48 0.26 0.52 0.22 0.67 0.44 0.41 0.21 0.64 0.29
$\sigma$-IG-F 0.69 0.35 0.88 0.52 0.60 0.42 0.43 0.40 0.56 0.29 0.49 0.24 0.56 0.20 0.66 0.44 0.38 0.22 0.64 0.29
Table 1: Class-wise AUCs for the pixel flipping plots for CIFAR-10. + and - denote the pixel insertion and pixel deletion metric respectively. The combination of the uncertainty estimation methods (Deep Ensemble (E), MC-Dropout (O), MC-DropConnect (C) and Flipout (F)), explanation methods (Guided BackPropagation (GBP) and Integrated Gradients (IG)), and type of explanation heatmap (mean ($\mu$) and standard deviation ($\sigma$)) are listed along the rows.
Deletion
$\mu$-GBP
$\sigma$-GBP
$\mu$-IG
$\sigma$-IG
Insertion
Deletion
Insertion
Figure 8: Pixel flipping curves for Deep Ensemble (rows 1-2) and MC-Dropout (rows 3-4) for FER+. The first and the third rows depict the deletion curves whereas the second and the fourth rows depict the insertion curves. From left to right, the columns show the following heatmaps and the explanation methods they correspond to: (i) $\mu$-GBP (ii) $\sigma$-GBP (iii) $\mu$-IG (iv) $\sigma$-IG.
Deletion
$\mu$-GBP
$\sigma$-GBP
$\mu$-IG
$\sigma$-IG
Insertion
Deletion
Insertion
Figure 9: Pixel flipping curves for MC-DropConnect (rows 1-2) and Flipout (rows 3-4) for FER+. The first and the third rows depict the deletion curves whereas the second and the fourth rows depict the insertion curves. From left to right, the columns show the following heatmaps and the explanation methods they correspond to: (i) $\mu$-GBP (ii) $\sigma$-GBP (iii) $\mu$-IG (iv) $\sigma$-IG.
contempt sadness neutral surprise anger happiness fear disgust
$\mu$-GBP-E 0.24 0.35 0.25 0.29 0.28 0.36 0.33 0.35 0.31 0.32 0.40 0.44 0.27 0.30 0.29 0.25
$\sigma$-GBP-E 0.25 0.34 0.25 0.28 0.28 0.35 0.33 0.34 0.32 0.32 0.40 0.43 0.27 0.29 0.31 0.24
$\mu$-IG-E 0.31 0.35 0.27 0.30 0.34 0.39 0.38 0.42 0.34 0.35 0.43 0.40 0.32 0.37 0.35 0.33
$\sigma$-IG-E 0.33 0.35 0.27 0.30 0.33 0.38 0.38 0.41 0.34 0.35 0.43 0.41 0.31 0.36 0.35 0.33
$\mu$-GBP-O 0.46 0.38 0.34 0.37 0.48 0.48 0.40 0.21 0.46 0.33 0.38 0.22 0.28 0.11 0.39 0.24
$\sigma$-GBP-O 0.45 0.36 0.37 0.33 0.48 0.43 0.39 0.21 0.47 0.31 0.41 0.22 0.31 0.18 0.33 0.21
$\mu$-IG-O 0.46 0.41 0.36 0.32 0.49 0.43 0.40 0.26 0.46 0.36 0.45 0.24 0.30 0.16 0.45 0.31
$\sigma$-IG-O 0.47 0.39 0.37 0.31 0.48 0.42 0.44 0.28 0.44 0.36 045. 0.25 0.31 0.20 0.44 0.29
$\mu$-GBP-C 0.21 0.29 0.22 0.25 0.23 0.27 0.21 0.23 0.21 0.25 0.27 0.34 0.16 0.19 0.16 0.23
$\sigma$-GBP-C 0.19 0.28 0.23 0.25 0.23 0.27 0.21 0.25 0.23 0.23 0.27 0.36 0.16 0.18 0.16 0.21
$\mu$-IG-C 0.28 0.22 0.30 0.24 0.31 0.26 0.32 0.24 0.26 0.23 0.39 0.28 0.27 0.24 0.31 0.20
$\sigma$-IG-C 0.29 0.25 0.29 0.23 0.30 0.26 0.34 0.24 0.26 0.22 0.37 0.28 0.28 0.26 0.26 0.21
$\mu$-GBP-F 0.27 0.32 0.31 0.29 0.32 0.33 0.46 0.27 0.27 0.25 0.32 0.24 0.39 0.28 0.28 0.19
$\sigma$-GBP-F 0.30 0.31 0.30 0.30 0.32 0.35 0.45 0.27 0.26 0.26 0.32 0.25 0.36 0.24 0.29 0.21
$\mu$-IG-F 0.30 0.31 0.35 0.33 0.40 0.36 0.48 0.33 0.30 0.26 0.44 0.22 0.39 0.33 0.29 0.32
$\sigma$-IG-F 0.38 0.32 0.35 0.36 0.40 0.39 0.48 0.35 0.29 0.29 0.47 0.23 0.39 0.32 0.31 0.32
Table 2: Class-wise AUCs for the pixel flipping plots for FER+. + and - denote the pixel insertion and pixel deletion metric respectively. The combination of the uncertainty estimation methods (Deep Ensemble (E), MC-Dropout (O), MC-DropConnect (C) and Flipout (F)), explanation methods (Guided BackPropagation (GBP) and Integrated Gradients (IG)), and type of explanation heatmap (mean ($\mu$) and standard deviation ($\sigma$)) are listed along the rows.
5.3 Explanation Uncertainty - Prediction vs Ground Truth Neurons
The presentation of the results is similar to that in Section 5.1 for the same input image. The results are shown in Figure 10.

### v2

@@ 1 Introduction
(a) Input
(b) $E_{\mu}$
(c) $E_{\sigma}$
Figure 1: Visualization of explanation with uncertainty on an image from FER+ dataset using Guided Backpropagation as explanation method and Deep Ensembles as uncertainty estimation method. (a) Input Image, (b) Explanation Mean, and (c) Explanation Uncertainty.
@@ 3.1 Explanation Uncertainty
$\mu(x)=T^{-1}\sum_{i}f_{\theta}(x),$ $\sigma^{2}(x)=T^{-1}\sum_{i}(f_{\theta}(x)-\mu(x))^{2}$ (7)
@@
where $\text{E}_{\sigma}(x)$ and $\text{E}_{\mu}(x)$ are the standard deviation and mean explanations respectively. The higher the coefficient of variation, the more the dissimilar the individual explanation heatmaps are. This metric can be helpful to quantify the confidence in the explanation of the network output for particular input features. Note that the explanation uncertainty is explicitly concerned with modeling uncertainty in the explanation and is unrelated to uncertainty explanation where the objective is to explain the uncertainty in a given prediction.
@@ 4 Experimental Setup
4.1 Image Classification Experiments
We test the proposed approach for the image classification task using the CIFAR-10 and the FER+ datasets. For this, we use the {Deep Ensemble, MC-Dropout, MC-DropConnect, Flipout} as uncertainty estimation methods and {Guided Backpropagation, Integrated Gradients} as the explanation methods. This provides us with a total of eight combinations of uncertainty estimation and explanation methods that we test. It should be noted that the LIME has not been used along with the image classification task as it is computationally expensive.
@@ 4.1 Image Classification Experiments
Pixel Deletion/Insertion. We focus on quantifying the quality of the generated explanations using our proposed approach. To this end, we identify the pixel deletion and the pixel insertion metric proposed by [10] in their work. The objective is to compute the changes in the predicted class score by the subsequent addition/removal of pixels to/from an image. We plot this change in the class score against the number of pixels added/deleted and compute the area under this curve (AUC) which is the pixel insertion/deletion metric. The addition/removal of the pixels to/from the input image is governed by the explanation heatmap associated with the image under consideration.
@@
Prediction vs Ground Truth Neurons. We propose modifying the computation of explanation uncertainty by varying the gradient computation step in our original approach. This modification is only tested with the FER+ dataset and is explained with the help of Figure 3.
Figure 3: Illustration describing an alternative approach to compute gradients and generate explanations. The original approach generates explanations by computing the gradient of the network prediction logit with respect to the input image whereas, in the modified variant, we compute the gradient of the ground truth logit with respect to the input image.
In our approach, the explanation is generated by calculating the gradient of the network output with respect to the input image. Mathematically, this is formulated as:
$E(x)=\frac{\partial(\text{Output}_{\text{Pred}})}{\partial(\text{Input}\ \text{Image})},\,E^{\prime}(x)=\frac{\partial(\text{Output}_{\text{GT}})}{\partial(\text{Input\ Image)}}$ (12)
where $\text{Output}_{\text{Pred}}$ is the same as network prediction from Figure 3. The explanation generated from this calculation helps us understand why the network predicted what it predicted.
However, now we modify the gradient computation step as can be seen in the Equation 12, where $\text{Output}_{\text{Gt}}$ is the activation of the unit corresponding to the ground truth (GT) as depicted in Figure 3. The explanation ($E^{\prime}(x)$) generated using this calculation helps to visualize the features/image regions that the network should have focused on for its prediction to match the ground truth. It is important to note that formulations listed in Equation 12 converge when the network prediction matches the ground truth for a given input.
@@
4.2 Regression Experiments
@@ 5 Experimental Results
Input Refer to caption
$E_{\mu}$ IG $E_{\sigma}$ IG CV IG $E_{\mu}$ GBP $E_{\sigma}$ GBP CV GBP
Ensemble Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-D Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MC-DC Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Visualization of the mean ($E_{\mu}$), the standard deviation ($E_{\sigma}$), and the coefficient of variation ($CV$) heatmaps generated using the combinations of uncertainty estimation methods (Deep Ensemble, MC-Dropout, MC-DropConnect and Flipout) and explanation methods (Guided Backpropagation and Integrated Gradients). The uncertainty estimation methods are organized in individual rows and the explanation methods are aligned along the columns with the columns 1-3 representing the $\mu$, $\sigma$, and the $CV$ heatmaps for IG and the columns 4-6 representing the same for GBP.
5.1 Image Classification Results
The explanation heatmaps for an image from the CIFAR-10 dataset are visualized in Figure 4. As previously described, the size of these heatmaps is the same as the size of the input image. In the subsequent discussions, the heatmaps will be referred to by their identifier namely the mean explanation heatmap as $\mu$, the standard deviation explanation heatmap as $\sigma$, and the coefficient of variation explanation heatmap as $CV$.
@@ 5.1 Image Classification Results
CIFAR10 FER+
Insert $\uparrow$ Delete $\downarrow$ Insert $\uparrow$ Delete $\downarrow$
$E_{\mu}$ $E_{\sigma}$ $E_{\mu}$ $E_{\sigma}$ $E_{\mu}$ $E_{\sigma}$ $E_{\mu}$ $E_{\sigma}$
GBP Ensemble 0.537 0.534 0.554 0.540 0.263 0.268 0.333 0.324
GBP MC-Dropout 0.551 0.541 0.320 0.323 0.354 0.347 0.293 0.281
GBP MC-DropConnect 0.582 0.571 0.368 0.354 0.186 0.187 0.256 0.253
GBP Flipout 0.586 0.557 0.301 0.299 0.291 0.289 0.271 0.274
IG Ensemble 0.547 0.542 0.606 0.571 0.304 0.304 0.364 0.361
IG MC-Dropout 0.582 0.571 0.368 0.345 0.374 0.378 0.311 0.301
IG MC-DropConnect 0.585 0.562 0.310 0.310 0.271 0.266 0.239 0.244
IG Flipout 0.598 0.589 0.334 0.337 0.328 0.341 0.308 0.323
Table 1: Area under the Insertion/Deletion curves, average across all classes, for CIFAR10 and FER+ datasets, computed separately for explanation mean $E_{\mu}$ and explanation uncertainty $E_{\sigma}$. The up/down arrows indicate direction of best metric value.
Similar analysis can be conducted for images from the FER+ dataset. The explanation distribution obtained for an image taken from FER+ are presented in the first four rows of Figure 5. It can be observed in the $\mu$ heatmaps that the GBP highlights the relevant facial features to a greater extent as compared to IG. This observation is supported by the $\sigma$ heatmaps of both these methods as well. The $\sigma$ and the $CV$ heatmaps generated by the GBP have low activation around the relevant pixel regions for all the uncertainty estimation methods. This implies that the explanations generated by the GBP have a higher agreement and are therefore highly confident. However, this is not the case for IG where the pixel activations in the $\mu$ heatmaps are scattered. This indicates a low confidence in the generated explanation.
@@
Pixel Deletion/Insertion. The results for pixel deletion/insertion are shown in Figure 9 (in appendix) and average AUC are shown in Figure 1. The tables containing the AUCs for these plots are provided in the Supplementary Materials. It should be noted that a higher AUC for pixel insertion and a lower AUC for pixel deletion is desired to ensure good-quality explanations. The pixel insertion and deletion curves (see Supplementary Material) have been plotted for batches of images representing all the classes in both datasets (CIFAR-10 and FER+). The order of pixel insertion/deletion is determined using the mean and the standard deviation explanation heatmaps of the corresponding images from these datasets.
It is evident from these results that for CIFAR-10, the best explanations are generally generated for the frog class and the worst explanations are generated for the dog class. These observations are supported by high AUC for pixel insertion and low AUC for pixel deletion for these classes. Overall across uncertainty methods and datasets, Flipout seems to perform the best, while Ensembles seems to perform the worse, specially in deletion curve for CIFAR10, but the variation between area under insertion/deletion curve is small.
@@
5.2 Prediction vs Ground Truth Neurons
The presentation of the results is similar to that in Section 5.1 for the same input image. The results are shown in the last four rows of Figure 5.